### v1

# From South to North: Leveraging Ground-Based LATs for Full-Sky CMB Delensing and Constraints on $r$

Abstract
Delensing—the process of mitigating the lensing-induced B-mode contamination in cosmic microwave background (CMB) observations—will be a pivotal challenge for next-generation CMB experiments seeking to detect primordial gravitational waves (PGWs) through B-mode polarization. This process requires an accurate lensing tracer, which can be obtained either through internal reconstruction from high-resolution CMB observations or from external large-scale structure (LSS) surveys. Ground-based large-aperture telescopes (LATs) are crucial for internal reconstruction, yet existing and planned facilities are confined to the southern hemisphere, limiting effective delensing to that region. In this work, we assess the impact of introducing a northern hemisphere LAT, assumed to be situated near AliCPT (hence termed Ali-like LAT, or LATN), on delensing performance and PGW detection, using simulations. Our baseline setup includes a space-based small-aperture mission (LiteBIRD-like, SAT) and a southern LAT (SO-like, LATS). External LSS tracers, which have been shown to play an important role in delensing before the availability of ultra-sensitive polarization data, are also considered. We find that southern-hemisphere internal delensing reduces the uncertainty in $r$ by $\sim$21compared to the no-delensing case. Adding LATN enables full-sky internal delensing, achieving a further $\sim$19reduction—comparable to that from including LSS tracers ($\sim$17%). Once LATN is included, the marginal benefit of LSS tracers drops to $\sim$10%. These results highlight the significant role of LATN in advancing delensing capabilities and improving PGW constraints.
1 Introduction
The Cosmic Microwave Background (CMB) radiation, a faint glow left over from the early universe, serves as a crucial window into the cosmos’ infancy. As one of the most important sources of information about the early universe, the CMB anisotropies provide a detailed snapshot of cosmic conditions just a few hundred thousand years after the Big Bang, offering critical insights into cosmology. In particular, the study of CMB B-mode polarization patterns has emerged as a powerful tool in the search for primordial gravitational waves (PGWs) [1, 2, 3, 4], which are predicted to be a direct consequence of inflation in the early universe. Detecting these modes would open a unique observational window into the physics of the very early universe, enabling us to probe fundamental questions about the origin of cosmic structures, the dynamics of inflation, and the fundamental nature of gravity.
The BICEP/Keck Array collaborations have recently placed constraints on PGWs [5, 6], parameterized by the tensor-to-scalar ratio $r$ (evaluated at a pivot scale of $0.05\,\text{Mpc}^{-1}$), obtaining an upper bound of $r<0.036$ at $2\sigma$ confidence for a fixed cosmology. Additionally, Ref. [7] combines BICEP/Keck Array data with $Planck$ PR4 and baryon acoustic oscillations (BAO), yielding a tighter constraint of $r<0.032$ at $2\sigma$ by simultaneously fitting cosmological parameters. Several ongoing and upcoming CMB experiments, including the BICEP Array [8], Simons Array [9], Simons Observatory (SO) [10], AliCPT [11], LiteBIRD [12], and CMB-S4 [13], aim to detect primordial B-mode polarization over the next decade, further advancing our ability to probe the early universe.
The precise measurement of the tensor-to-scalar ratio $r$ through cosmic microwave background (CMB) B-mode polarization is hindered by various sources of contamination. One of the major challenges arises from gravitational lensing by large-scale structure (LSS), which introduces a significant and unavoidable source of confusion. This lensing converts primordial E-modes into secondary B-modes, thereby mimicking the signal from primordial gravitational waves and obscuring the detection of a true $r$. When $r\lesssim 0.03$[12], lensing-induced B-modes dominate the signal, necessitating their removal through delensing to recover the primordial component. Next-generation CMB experiments aim to suppress this lensing contamination via high-resolution lensing reconstruction and cross-correlation with external LSS tracers, thus tightening constraints on $r$ and enabling more robust tests of inflationary cosmology [14, 15, 16, 13]. Polarized Galactic emission—mainly from thermal dust and synchrotron radiation by relativistic electrons—is another major contaminant in CMB observations [17]. To mitigate this, component separation methods such as Commander[18, 19], NILC[20], SEVEM[21], and SMICA[22] have been developed to isolate foregrounds based on their distinct spectral signatures.
The removal of lensing-induced contamination from CMB anisotropy maps has been a major focus in recent literature [23, 24, 25, 26]. Accurate delensing requires prior knowledge of the lensing potential $\phi$, which can be obtained through various methods in practice. One approach involves internal reconstruction using quadratic estimators based on CMB two-point statistics [27], while Bayesian iterative methods [28, 29] provide a statistically optimal alternative. Additionally, external mass tracers, such as the Cosmic Infrared Background and high-redshift galaxy surveys, act as effective proxies for lensing and have shown significant improvements, particularly in the absence of future polarization surveys with deeper sensitivities [30, 31, 32]. For the internal lensing reconstruction, high-resolution ($-\ell$) observation with low instrumental noise is necessary to reduce reconstruction uncertainty. Ground-based telescopes are exceptionally well-suited for this task, as their large apertures—readily achievable in ground-based installations—enable precise measurements of small-scale CMB fluctuations. Although atmospheric noise typically degrades large-scale sensitivity, its impact on lensing reconstruction is limited if properly accounted for. Furthermore, foreground contamination across multiple frequency channels (27–280,GHz) can be effectively mitigated using the Needlet Internal Linear Combination (NILC) method.
Satellite-based small-aperture missions such as LiteBIRD enable full-sky measurements of large-scale CMB B-mode polarization, offering significant advantages over ground-based SATs due to their wide sky coverage and broad frequency range. The extensive sky coverage helps reduce sample variance in measuring the tensor-to-scalar ratio $r$, while the richer frequency information enhances component separation between the CMB and foregrounds. However, to fully realize the potential of such missions in constraining $r$, delensing is essential. Ref.[16] reports a 20% reduction in the uncertainty of $r$ using a multi-tracer delensing approach, despite residual foreground contamination. This study employs mock internal lensing reconstruction based on CMB-S4 LAT data (primarily from the Southern Hemisphere) and incorporates external LSS tracers, such as the Cosmic Infrared Background (CIB), Euclid[33], and LSST [34], to enhance the mass distribution estimate and demonstrate the effectiveness of multi-tracer delensing for future full-sky surveys.
For future ultra-low-noise observations [29], internal lensing reconstruction is expected to provide more accurate lensing estimates. In this context, achieving a full-sky internal reconstruction of the lensing potential using ground-based large-aperture telescopes (LATs) is crucial for enabling full-sky delensing and tightening constraints on the tensor-to-scalar ratio $r$. In this work, to emphasize the importance of full-sky internal delensing, we focus on a near-future achievable configuration, using a southern LAT (SO-like, denoted as LATS) as the baseline, and assess the additional contribution of a northern LAT, assumed to be situated near AliCPT[35] (hence termed Ali-like LAT, or LATN), to full-sky primordial gravitational wave (PGW) detection through delensing. Our simulations combine polarization data from a hypothetical satellite mission (SAT)—featuring ultra-low instrumental noise, no atmospheric contamination, and wide frequency coverage (40–402 GHz across 15 channels)—with data from both LATs. Foreground removal is performed using the NILC method. We demonstrate that, even with a near-future LAT configurations, the inclusion of LATN, enabling full-sky internal delensing, yields a substantial improvement in constraints on $r$—comparable to that achieved through the use of external large-scale structure tracers.
This paper is organized as follows. In Section 2, we introduce the delensing methodology along with the associated debiasing formalism. Section 3 outlines the complete analysis pipeline, including map simulation, foreground cleaning, map co-addition, lensing reconstruction, CMB B-mode delensing, and parameter estimation. In Section 4, we present the performance of each step in the pipeline, culminating in the final parameter constraints. We conclude the paper in Section 5.
2 Large-scale B-mode Delensing
As cosmic microwave background (CMB) photons travel from the last-scattering surface to Earth, they encounter large-scale structures (LSS), which induce a cumulative gravitational lensing effect. This effect displaces the primordial CMB photons, causing them to reach us from directions that deviate from their original line of sight (see [36] for a comprehensive review). Mathematically, this process can be described as a remapping of the CMB temperature and polarization anisotropies:
$\tilde{\Theta}(\hat{\mathbf{n}})$ $=\Theta(\hat{\mathbf{n}}+\mathbf{d}(\hat{\mathbf{n}}))=\Theta(\hat{\mathbf{n}})+\nabla_{i}\Theta(\hat{\mathbf{n}})\mathbf{d}^{i}(\hat{\mathbf{n}})+\mathcal{O}(\phi^{2}),$ (2.1)
$[\tilde{Q}\pm i\tilde{U}](\hat{\mathbf{n}})$ $=[Q\pm iU](\hat{\mathbf{n}}+\mathbf{d}(\hat{\mathbf{n}}))=[Q\pm iU](\hat{\mathbf{n}})+\nabla_{i}[Q\pm iU](\hat{\mathbf{n}})\mathbf{d}^{i}(\hat{\mathbf{n}})+\mathcal{O}(\phi^{2}),$ (2.2)
where tildes indicate lensed quantities and $\mathbf{d}$ is the deflection angle, given by the gradient of the lensing potential $\mathbf{d}(\hat{\mathbf{n}})=\nabla\phi(\hat{\mathbf{n}})$ (here we ignore the curl modes) [37], where the lensing potential is the projection of gravitational potential along line-of-sight [27]:
[…]
where $\eta$ is the conformal time, $\eta_{s}$ is the epoch of last scattering and $\chi$ is the angular diameter distance in comoving coordinates.
One interesting aspect for polarization is that in harmonic space, the remapping will convert part of the E-modes into B-modes, known as lensing B-modes. At linear order in $\phi$, these are given by (neglecting the tensor B-modes) [14]:
${B}_{\ell m}^{\text{lens}}$ $=i\sum_{\ell^{\prime}m^{\prime}}\sum_{LM}\begin{pmatrix}\ell&\ell^{\prime}&L\\ m&m^{\prime}&M\end{pmatrix}p^{-}F^{(2)}_{\ell L\ell^{\prime}}E^{*}_{\ell^{\prime}m^{\prime}}\kappa^{*}_{LM},$ (2.3)
where $p^{+}(p^{-})$ is unity if $\ell+L+\ell^{\prime}$ is even (odd) and zero otherwise. Lensing induced mode coupling for $s$ fields can be expressed as:
$F^{(s)}_{lLl^{\prime}}=\frac{2}{L(L+1)}\left[l^{\prime}(l^{\prime}+1)+L(L+1)-l(l+1)\right]$ (2.4)
×$\times\sqrt{\frac{(2l+1)(2l^{\prime}+1)(2L+1)}{16\pi}}\begin{pmatrix}\ell&\ell^{\prime}&0\\ -s&s^{\prime}&0\end{pmatrix}$
As instrumental sensitivity improves, lensing-induced B-modes have become an increasingly significant source of contamination in the observation of primordial gravitational waves (PGWs). Considerable attention has been devoted to mitigating lensing B-modes from observations, and various methods have been developed to achieve this goal (see, e.g., [23, 24, 25, 26]). This process, known as CMB B-mode delensing, aims to reduce the lensing contribution to enhance the detectability of primordial signals. Here, we briefly describe two delensing methods used in this work.
2.1 Delensing Method Description
Gradient-Order Template Method
A straightforward approach to mitigating lensing B-modes is to construct a lensing B-mode template at gradient order, which serves as a good approximation for lensing B-modes on large scales [38, 26]. It is convenient to work in real space, drawing inspiration from Eq. (2.2):
$\nabla_{i}[Q\pm iU](\hat{\mathbf{n}})\mathbf{d}^{i}(\hat{\mathbf{n}}).$ (2.5)
In practice, the gradient of a field can be efficiently computed using the ladder operators [27]:
$\begin{gathered}D_{i}[{{}_{s}}f(\hat{\mathbf{n}})]=-\frac{1}{\sqrt{2}}\left\{\sharp{{}_{s}}f(\hat{\mathbf{n}})\bar{m}+\flat{{}_{s}}f(\hat{\mathbf{n}})m\right\},\end{gathered}$ (2.6)
where the covariant derivative $\nabla_{i}$ operating on the $s$ field piece of a tensor is equivalent to a gradient operation $D_{i}$ on its $s$ weighted representation, $\sharp$ and $\flat$ denote the ladder operators, and the complex-conjugated vectors $\bar{m}$ and $m$ serve as the basis.
A crucial point to note is that the $Q$ and $U$ fields used to construct the template must exclude the B-modes. Otherwise, the lensing operation would be applied again, introducing redundant terms. The resulting template can then be subtracted from the observed map (map-level delensing) or treated as a pseudo-channel in the likelihood analysis when constraining parameters (cross-spectral method delensing), as discussed in Section 4.3.
Inverse-lensing Method
Another method is more intuitive, by reversing the lensing effect through remapping the observed photons back to their original positions, we can achieve a more accurate and optimal approach for delensing. The inverse deflection angle is well-defined because the points are remapped onto themselves after being deflected back and forth, as discussed in [24]:
$\hat{\mathbf{n}}+\mathbf{d^{inv}}(\hat{\mathbf{n}})+\mathbf{d}(\hat{\mathbf{n}}+\mathbf{d^{inv}}(\hat{\mathbf{n}}))=\hat{\mathbf{n}},$ (2.7)
where the primary CMB fields can be recovered from the lensed field by remapping the latter using the inverse deflection angle $\mathbf{d^{inv}}(\hat{\mathbf{n}})$:
$\tilde{X}(\hat{\mathbf{n}})=X(\hat{\mathbf{n}}+\mathbf{d}(\hat{\mathbf{n}}))\Leftrightarrow X(\hat{\mathbf{n}})=\tilde{X}(\hat{\mathbf{n}}+\mathbf{d^{inv}}(\hat{\mathbf{n}})).$ (2.8)
In practice, at low-order approximation, the inverse deflection angle is approximately the negative of the deflection angle. A more accurate estimate can be obtained by solving Eq. (2.7) using a Newton-Raphson scheme, where the inverse deflection angle $\mathbf{d^{\text{inv}}}(\hat{\mathbf{n}})$ is iteratively computed, as described in [25]. The lensing B-mode template is then constructed by subtracting the inverse-lensed B-mode map from the observed data.
2.2 Bias Analysis
Although the delensing procedure mitigates part of the lensing effect in observations, it also introduces biases in the measurement of $r$. These biases arise from inherent correlations between the lensed CMB, noise and foreground residuals, and the lensing potential. Such correlations manifest across different statistical orders, including the two-point correlation (angular power spectrum), three-point correlation (bispectrum), and four-point correlation (trispectrum). Below we do not explicitly derive each analytical term. Instead, we present the analysis in a way that facilitates understanding their origins and estimating their magnitudes.
The NILC cleaned E-modes, B-modes and the estimated lensing potential with Wiener filters (for a detailed description, see Section 3.5), are as follows:
$E=\mathcal{W}^{E}(E^{\text{lens}}+E^{\text{res,NILC}}),$ (2.9)
$B=\mathcal{W}^{E}(B^{\text{lens}}+B^{\text{res,NILC}}),$
$\hat{\phi}=\mathcal{W}^{\phi}(\phi+\phi^{\text{noise}}),$
where $\mathcal{W}^{E}$ and $\mathcal{W}^{\phi}$ are the Wiener filters applied to CMB polarization observation and reconstructed lensing potential, respectively.
2.2.1 Bias Analysis of the Gradient-Order Template Method
Starting from the expression above, the lensing B-mode template can be written as:
$B^{\text{temp}}$ $=\mathcal{B}^{(1)}[\mathcal{W}^{E}(E^{\text{lens}}+E^{\text{res,NILC}})\ast\mathcal{W}^{\phi}(\phi+\phi^{\text{noise}})]$ (2.10)
=$&\mathcal{B}^{(1)}[\mathcal{W}^{E}E^{\text{lens}}\ast\mathcal{W}^{\phi}\phi]+\left\{\mathcal{B}^{(1)}[\mathcal{W}^{E}E^{\text{res,NILC}}\ast\mathcal{W}^{\phi}\phi]+\mathcal{B}^{(1)}[\mathcal{W}^{E}(E^{\text{lens}}+E^{\text{res,NILC}})\ast\mathcal{W}^{\phi}\phi^{\text{noise}}]\right\}\end{aligned}$
$=B^{\text{temp}}_{S}+B^{\text{temp}}_{N},$
where the first term represents the signal component of the template, while the remaining terms correspond to noise, foreground residuals, and lensing reconstruction noise. Notably, all of these contributions are suppressed by the Wiener filter.
The auto-power spectrum of the lensing template from Eq. (2.10) is then:
$C^{\text{temp}}=C_{S}^{\text{temp}}+C_{N}^{\text{temp}}+2\langle B^{\text{temp}}_{S}B^{\text{temp}}_{N}\rangle.$ (2.11)
The cross-power spectrum between $B^{\text{temp}}=\mathcal{T}^{-1}B^{\text{lens}}+B_{N}^{\text{temp}}$ and $B^{\text{obs}}=B^{\text{lens}}+B^{\text{res,NILC}}$ is:
$C^{\text{cross}}=\mathcal{T}^{-1}C^{\text{lens}}+\langle B^{\text{lens}}B_{N}^{\text{temp}}\rangle+\langle B^{\text{res,NILC}}B^{\text{temp}}\rangle,$ (2.12)
where $\mathcal{T}$ is the transfer function that compensates for signal attenuation due to the Wiener filter.
In practice, we find that most of the bias terms are realization-independent. Therefore, to estimate these significant bias terms, we can either compute them directly through simulations or extract them by taking the difference between the baseline result and the results obtained using the signal-only simulations.
2.2.2 Bias Analysis of the Inverse-Lensing Method
Similarly, we start from the remapped B-modes:
$B^{\text{del}}$ $=\mathcal{B}[\mathcal{W}^{E}(B^{\text{lens}}+B^{\text{res,NILC}})\star\mathbf{d^{inv}}]$ (2.13)
$\approx\mathcal{B}[\mathcal{W}^{E}(B^{\text{lens}}+B^{\text{res,NILC}})\star(\mathbf{d^{inv}_{\phi}}+\mathbf{d^{inv}_{\text{noise}}})]$
$\approx\left\{\mathcal{B}[\mathcal{W}^{E}B^{\text{lens}}\star\mathbf{d^{inv}_{\phi}}]\right\}+\left\{\mathcal{B}[\mathcal{W}^{E}B^{\text{lens}}\star\mathbf{d^{inv}_{\text{noise}}}]+\mathcal{B}[\mathcal{W}^{E}B^{\text{res,NILC}}\star\mathbf{d^{inv}}]-\mathcal{W}^{E}B^{\text{lens}}\right\}$
$=B^{\text{del}}_{S}+B^{\text{del}}_{N},$
The lensing template is obtained by subtracting the delensed B-mode from the filtered observed B-modes:
$B^{\text{temp}}$ $=\mathcal{W}^{E}B^{obs}-B^{\text{del}}$ (2.14)
$=\mathcal{W}^{E}[B^{\text{lens}}+B^{\text{res,NILC}}]-[B^{\text{del}}_{S}+B^{\text{del}}_{N}]$
$\approx[\mathcal{W}^{E}B^{\text{lens}}-B^{\text{del}}_{S}]+[\mathcal{W}^{E}B^{\text{res,NILC}}-B^{\text{del}}_{N}]$
$=B^{\text{temp}}_{S}+B^{\text{df}},$
where the same filter is applied to the observed B-modes to ensure consistency with the delensed B-modes, which experience signal attenuation due to the Wiener filter.
Next, we calculate the auto-power spectrum of the lensing template, given by $B^{\text{temp}}=\mathcal{T}^{-1}B^{\text{lens}}+B^{\text{df}}$:
$C^{\text{temp}}$ $=\mathcal{T}^{-2}C^{\text{lens}}+\langle B^{\text{df}}B^{\text{df}}\rangle+2\mathcal{T}^{-1}\langle B^{\text{lens}}B^{\text{df}}\rangle,$ (2.15)
where $\mathcal{T}$ represents the transfer function used to compensate for signal attenuation caused by the Wiener filter.
For the cross-power spectrum with the observed B-modes, $B^{\text{obs}}=B^{\text{lens}}+B^{\text{res,NILC}}$, we have:
$C^{\text{cross}}$ $=\mathcal{T}^{-1}C^{\text{lens}}+\langle B^{\text{lens}}B^{\text{df}}\rangle+\langle B^{\text{res,NILC}}B^{\text{temp}}\rangle.$ (2.16)
These bias terms essentially correspond to those discussed in the previous section, with the exception of some minor high-order effects that depend on the accuracy of the inverse deflection angle $\mathbf{d^{inv}}(\hat{\mathbf{n}})$. Therefore, they can be estimated in a similar manner to the latter.
3 Step-by-Step Delensing Process: Techniques and Implementation
In this section, we present the complete delensing pipeline implemented in our simulation-based analysis. The workflow consists of the following key stages: We begin with an overview of the simulated datasets used throughout this work (Section 3.1), followed by a description of the component separation procedure implemented via NILC (Section 3.2). We then detail the implementation of lensing reconstruction—both internal and external (Section 3.4). These reconstructions are then used to construct large-scale CMB lensing B-mode templates (Section 3.5). Finally, we perform a likelihood analysis using Markov Chain Monte Carlo (MCMC) methods to derive constraints on cosmological parameters (Section 3.6). A schematic representation of the full workflow is provided in Fig. 1, highlighting the key steps and their interdependencies.
3.1 Data Simulation
We employ simulated data to evaluate the feasibility of the delensing pipeline. Specifically, we simulate future CMB observations from various experiments, including a large-aperture, ground-based 666-m telescope in the Northern Hemisphere (LATN), a large-aperture, ground-based 666-m telescope in the Southern Hemisphere (LATS), and a small-aperture 111-m satellite telescope (SAT). We then assess the effectiveness of different delensing methods in reducing the uncertainty in $r$ as shown in Section 4.3.
Figure 1: The flowchart of the entire analysis pipeline is presented for clarity. The workflow is broadly divided into five main components: component separation using NILC, map combination, lensing reconstruction, construction of the lensing $B$-mode template, and parameter constraint estimation. While the aim is not to detail every operational step, the flowchart highlights the logical structure and data dependencies throughout the process. For reference, the resulting reduction of uncertainties in $r$ are also summarized alongside the corresponding configurations. Note that the arrow from $\kappa$ to the lensing template (LT) construction indicates a computational pathway only, and does not imply that the lensing B-modes are physically induced by the convergence field.
Figure 2: Survey windows for each observation in Galactic coordinates used in the analysis. For the coadded case, we combine two ground-based LATs with the satellite SAT. The yellow region represents the overlap between SAT and one of the LATs, while the brown region indicates the common region covered by all three observations.
The survey windows for these observations are shown in Fig. 2. The two ground-based LATs have sky coverages of approximately 38% and 43% after masking the Galactic disk, respectively, with an overlap of about 20% (depicted in brown in the bottom-right panel of Fig. 2). The satellite SAT covers 70% of the sky after masking the Galactic disk, with an overlap of approximately 61% with the two LATs, represented by the combined brown and yellow regions in the bottom-right panel of Fig. 2
The input unlensed CMB maps are Gaussian realizations generated from a specific power spectrum obtained using the Boltzmann code CAMB [39], based on the Planck 2018 best-fit cosmological parameters [40] with a tensor-to-scalar ratio $r=0$. The Gaussian realization of the lensing potential maps is also generated from the fiducial lensing potential power spectrum. We then utilize the Lenspyx [41, 42] package, which implements an algorithm to distort the primordial signal based on the lensing potential map generated above.
We include both Galactic and extragalactic foregrounds in our simulation using PySM3 [43, 44]. PySM3 generates full-sky simulations of Galactic emissions in intensity and polarization based on realistic space observations (e.g., Planck, WMAP). The modeled diffuse Galactic emissions include thermal dust (d9 model), synchrotron (s4 model), spinning dust (a1 model), and free-free emission (f1 model). The extragalactic emissions incorporated in our simulations include the Sunyaev-Zel’dovich (SZ) effect (ksz1 and tsz1 models) and the cosmic infrared background (cib1 model).
Since the foreground templates described above are derived from real observational data, only a single simulation is available. Consequently, our analysis does not account for uncertainties in foreground modeling. However, we apply the NILC method to clean the observations (see Section 3.2), and the foreground residuals are expected to be largely suppressed. Therefore, we anticipate that our results are not significantly overestimated.
For the noise simulation of ground-based LATs, we assume that both telescopes share the same experimental configuration and observational conditions, differing only in their survey windows. We adopt the noise model described in [45], which accounts for both detector white noise and a $1/f$ component. The total noise power spectrum is given by:
$N_{\ell}=N_{\text{red}}\left(\frac{\ell}{\ell_{\text{knee}}}\right)^{\alpha_{\text{knee}}}+N_{\text{white}},$ (3.1)
[…]
Table 1: The experimental parameters of the LAT experiments. $\sigma_{\text{white}}$ give white noise levels for temperature, with polarization noise $\sqrt{2}$ higher as both $Q$ and $U$ Stokes parameters. Both the LATN and LATS configurations adopt the ”goal” specifications define for SO surveys [45].
LAT Temperature LAT Polarization Freq. [GHz] FWHM [arcmin] $\sigma_{\text{white}}$ [$\mu$K-arcmin] $N_{\text{red}}\cdot t$ [$\mu$K2-s] $\ell_{\text{knee}}$ $\alpha_{\text{knee}}$ $N_{\text{red}}$ [$\mu$K2] $\ell_{\text{knee}}$ $\alpha_{\text{knee}}$ 27 7.4 52 100 1000 -3.5 $N_{\text{white,pol.}}$ 700 -1.4 39 5.1 27 39 1000 -3.5 $N_{\text{white,pol.}}$ 700 -1.4 93 2.2 5.8 230 1000 -3.5 $N_{\text{white,pol.}}$ 700 -1.4 145 1.4 6.3 1500 1000 -3.5 $N_{\text{white,pol.}}$ 700 -1.4 225 1.0 15 17000 1000 -3.5 $N_{\text{white,pol.}}$ 700 -1.4 280 0.9 37 31000 1000 -3.5 $N_{\text{white,pol.}}$ 700 -1.4
Table 2: Parameters of the SAT experiment. For the simplicity of simulation, the values shown are approximate estimates based on the science goals of LiteBIRD; for an updated and precise description of the instrumentation, see [46, 12].
Freq. [GHz] FWHM [arcmin]
Noise level [$\mu$K-arcmin]
40 70.5 26.5
50 58.5 23.7
60 51.1 15.1
68 41.6 14.1
78 36.9 11.0
89 33.0 8.7
100 30.2 7.3
119 26.3 5.4
140 23.7 5.1
166 28.9 3.9
195 28.0 5.0
235 24.7 7.6
280 22.5 9.8
337 20.9 15.5
402 17.9 33.6
(a) LAT temperature
(b) LAT polarization
Figure 3: The total noise power spectra (beam-corrected) for the ground-based LATs.
The noise for the SAT is homogeneous, generated through Gaussian random sampling on the map domain with noise levels specified in Table 2.
Finally, the observed maps are derived by summing the lensed CMB maps, foreground maps, and noise maps.
3.2 Needlet Internal Linear Combination
Component separation using the ILC (Internal Linear Combination) method [20] has been widely applied in CMB observations. In this work, we perform component separation using the NILC (Needlet Internal Linear Combination) method independently for each experiment. The expansion of sky maps in spherical needlets allows for the localization of statistical properties in both pixel space and harmonic space, thereby enhancing the ability to account for variations in foreground intensity across the sky and the dominance of instrumental noise at small scales. Under the assumption that each component of the observed maps is independent, the observed field can be expressed as
$X_{\ell m}^{\mathrm{obs},c}=b_{\ell}^{c}X_{\ell m}^{\mathrm{CMB}}+b_{\ell}^{c}X_{\ell m}^{\mathrm{FG},c}+X_{\ell m}^{\mathrm{N},c},$ (3.2)
where $X\in\{\Theta,E,B\}$ represents the harmonic coefficient of a field, $b_{\ell}^{c}$ is the beam function, and $c$ denotes the frequency channel. We then perform a filter $h_{\ell}^{j}$ in harmonic space to extract observation at the very scale $j$,
$X_{\ell m}^{c,j}=h_{\ell}^{j}X_{\ell m}^{c},$ (3.3)
this corresponds to expanding the observed harmonic coefficient with spherical needlets at scale $j$, pixel $k$:
$\beta_{jk}^{X,c}=\beta_{jk}^{\mathrm{CMB}}+\beta_{jk}^{\mathrm{FG},c}+\beta_{jk}^{\mathrm{N},c},$ (3.4)
with corresponding spherical needlets given by:
$\beta_{jk}^{\mathrm{CMB}}$ $=\sqrt{\lambda_{j}}\sum_{\ell m}h_{\ell}^{j}b_{\ell}X_{\ell m}^{\mathrm{CMB}}Y_{\ell m}(\xi_{j}),$ (3.5)
$\beta_{jk}^{\mathrm{FG},c}$ $=\sqrt{\lambda_{j}}\sum_{\ell m}h_{\ell}^{j}b_{\ell}^{c}X_{\ell m}^{\mathrm{FG},c}Y_{\ell m}(\xi_{j}),$
$\beta_{jk}^{\mathrm{N},c}$ $=\sqrt{\lambda_{j}}\sum_{\ell m}h_{\ell}^{j}\frac{b_{\ell}}{b_{\ell}^{c}}X_{\ell m}^{\mathrm{N},c}Y_{\ell m}(\xi_{j}).$
Since all the observations actually measure signal $bX^{\mathrm{CMB}}$ with some error $bX^{\mathrm{FG}}+X^{\mathrm{N}}$, consists in averaging all these measurements giving a specific weight $w_{i}$ to each of them, we naturally questing for a solution of the form in needlet space:
$\beta_{jk}^{\mathrm{NILC}}=\sum_{c=1}^{n_{c}}w_{jk}^{c}\beta_{jk}^{X,c},$ (3.6)
the weight $w_{jk}^{c}$ is determined by minimizing the variance of the reconstructed CMB, subject to the unbiasedness constraint ($\sum_{c=1}^{n_{c}}w_{jk}^{c}=1$). Using the method of Lagrange multipliers to enforce the constraint, the solution is given by:
$\hat{\omega}_{jk}^{c}=\frac{\sum_{c^{\prime}}\left[\hat{\mathbf{R}}_{jk}^{-1}\right]^{cc^{\prime}}a^{c^{\prime}}}{\sum_{c}\sum_{c^{\prime}}a^{c}\left[\hat{\mathbf{R}}_{jk}^{-1}\right]^{cc^{\prime}}a^{c^{\prime}}},$ (3.7)
where $\left[\hat{\mathbf{R}}_{jk}^{-1}\right]^{cc^{\prime}}$ is the inverse of the cross frequency covariance matrix:
$\hat{\mathbf{R}}_{jk}^{cc^{\prime}}=\frac{1}{n_{k}}\sum_{k^{\prime}}\omega_{j}(k,k^{\prime})\beta_{jk}^{c}\beta_{jk}^{c^{\prime}},$ (3.8)
where $\omega_{j}(k,k^{\prime})$ selects the domain of $n_{k}$ pixels around the $k$-th pixel over which we perform our averaging to estimate the covariance. Finally, the cleaned CMB map is given by:
$\hat{X}_{\text{NILC}}(\hat{\mathbf{n}})=\sum_{\ell m}\sum_{jk}\beta_{jk}^{\mathrm{NILC}}\sqrt{\frac{4\pi}{\lambda_{j}}}h_{\ell}^{j}Y_{\ell m}(\xi_{j})Y_{\ell m}(\hat{\mathbf{n}}).$ (3.9)
The harmonic transformations mentioned therein can be executed using healpy / HEALPix package  [^1] https://healpix.sourceforge.io/ [47, 48]. A well-tested Python package to implement the NILC pipeline, is available at openilc [^2] https://github.com/dreamthreebs/openilc.
For the two ground-based LATs, we use beam-deconvolved maps at six frequencies as inputs, as detailed in Table 1. For the satellite-based SAT, maps at fifteen frequencies are employed, as outlined in Table 2. Component separation is performed independently for the observed $\Theta$, $E$, and $B$ maps. Additionally, E-to-B leakage is corrected during the transformation from Stokes parameters $Q$ and $U$ to $E$ and $B$ using the ”E-mode recycling” method introduced by [49]. The final outputs consist of cleaned $\Theta$, $E$, and $B$ maps for each experiment, denoted as $X^{\text{NILC,obs},i}$, where $i$ represents the respective experiment.
Furthermore, we linearly combine the corresponding noise and foreground maps (i.e., nulling the signal in the observed maps) for each experiment using the weights obtained above. The resulting output maps, $X^{\text{NILC,null},i}$, are used exclusively for inverse-weight calculations in harmonic space for map combination and for estimating the effective white noise level (see Section 3.4).
3.3 Map combination
We combine the NILC-cleaned maps from each experiment using inverse-variance weighting in harmonic space. The inverse-variance combination is given by:
$\chi_{\ell m}=\sum_{i}\omega_{i,\ell}\chi_{i,\ell m},$ (3.10)
$\omega_{i,\ell}=\frac{N_{i,\ell}^{-1}}{\sum_{i}N_{i,\ell}^{-1}},$
where $\omega_{i,\ell}$ represents the weight of the $i$-th experiment, $N_{i,\ell}$ is the power spectrum of the noise and residual foregrounds from the $i$-th experiment, denoted as $X^{\text{NILC,null},i}$, $\chi_{i,\ell m}$ corresponds to the harmonic coefficients of the fields $\Theta,E,B$ for the $i$-th experiment after foreground cleaning, and $\chi_{\ell m}$ represents the combined harmonic coefficients.
It is important to note that inverse-variance weighting can only be applied to fields that cover the same sky regions and share the same multipole ranges; otherwise, they will be directly mapped onto a null sky map. For instance, consider two observed maps from different experiments, $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$, as illustrated in Fig. 4. Due to differences in experimental configurations—such as aperture size, location, atmospheric conditions, and scan strategy—their respective sky coverages differ. We denote the exclusive sky regions of these experiments as $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, while their overlapping region is denoted as $\mathcal{C}$. Similarly, their multipole coverage also differs. We define these ranges as $\mathcal{L}_{1}=\mathcal{B}_{1}+\mathcal{D}$ and $\mathcal{L}_{2}=\mathcal{B}_{2}+\mathcal{D}$, where $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ represent the experiment-specific multipole ranges, and $\mathcal{D}$ denotes the common multipole range.
The combination strategy follows these steps:
  1. 1.
    Identify the sky regions: First, determine the overlapping region ($\mathcal{C}$) and the exclusive regions for each experiment ($\mathcal{S}_{1}$ and $\mathcal{S}_{2}$) using the corresponding masks shown in Fig. 2. These masks have been apodized using the ”Smooth” function (effectively a Gaussian window) implemented in NaMaster[50], with an apodization scale of $6^{\prime}$ to reduce artifacts caused by sharp mask edges.
  2. 2.
    Identify the multipole ranges: Next, determine the common multipole range ($\mathcal{D}$) and the experiment-specific multipole ranges ($\mathcal{B}_{1}$ and $\mathcal{B}_{2}$).
  3. 3.
    Decompose information by region and scale: The maps corresponding to each region and multipole range are extracted using a spherical harmonic transform, with the masks obtained in the previous step applied. This results in the following harmonic coefficients:
    $a_{\mathcal{B}_{1}+\mathcal{D}}^{\mathcal{S}_{1}},\quad b_{\mathcal{B}_{2}+\mathcal{D}}^{\mathcal{S}_{2}},\quad a_{\mathcal{D}}^{\mathcal{C}},\quad b_{\mathcal{D}}^{\mathcal{C}},\quad a_{\mathcal{B}_{1}}^{\mathcal{C}},\quad b_{\mathcal{B}_{2}}^{\mathcal{C}}.$
    Here, $m_{\mathcal{L}}^{\mathcal{S}}$ represents the harmonic coefficients of experiment $m$ ($a$ for $\mathcal{M}_{1}$, $b$ for $\mathcal{M}_{2}$) in region $\mathcal{S}$ and at multipole range $\mathcal{L}$.
  4. 4.
    Combine the maps in the pixel domain, region by region:
    1. (a)
      For region $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, the output is simply $a_{\mathcal{B}_{1}+\mathcal{D}}^{\mathcal{S}_{1}}$ and $b_{\mathcal{B}_{2}+\mathcal{D}}^{\mathcal{S}_{2}}$, respectively. It is worth noting that these cases require no spherical harmonic transformation in practice; a simple reprojection at the map level suffices.
    2. (b)
      For the common region $\mathcal{C}$, perform a weighted combination as given by Eq. (3.10) on $a_{\mathcal{D}}^{\mathcal{C}}$ and $b_{\mathcal{D}}^{\mathcal{C}}$. The final output for this region is:
      $\omega_{1}a_{\mathcal{D}}^{\mathcal{C}}+\omega_{2}b_{\mathcal{D}}^{\mathcal{C}}+a_{\mathcal{B}_{1}}^{\mathcal{C}}+b_{\mathcal{B}_{2}}^{\mathcal{C}}.$
(a) Observation regions of two example experiments.
(b) Multipole range coverage of two example experiments.
Figure 4: Illustration of the sky coverage and multipole range for two example experiments used in the map combination. Panel (a) shows the sky regions observed by each experiment, while panel (b) represents the corresponding multipole range coverage.
As a result, the combined map is inverse-variance weighted only in the overlapping regions at the common scales, while preserving the remaining information. This approach is reasonable, as it fully utilizes the available information for weighted combination—i.e., no additional information is accessible for weighting in the exclusive regions or at the private scales.
We emphasize that this work employs the diagonal filtering approach, despite its known sub-optimal performance for cut-sky analyses when compared to the more sophisticated methodology described in [51]. This decision is based on two considerations. First, we have tested the alternative method and found it to be computationally demanding. Second, we do not expect a significant improvement in delensing efficiency from this adjustment. As described by [52], the key factor determining delensing efficiency is the accuracy of the lensing estimate, rather than contamination in the CMB maps undergoing delensing.
In this paper, we first combine the NILC-cleaned maps from the two LATs (hereafter denoted as LATN+LATS), both of which cover a multipole range of (0,6143)(0,6143)( 0 , 6143 ), over the sky regions shown in Fig. 2. The outputs are the combined $\Theta$, $E$, and $B$ maps. These maps serve two purposes. First, they are used for internal lensing reconstruction. Second, they are further combined with the NILC-cleaned maps from the SAT (hereafter denoted as LATN+LATS+SAT), where the combined LAT covers the multipole range (0,6143)(0,6143)( 0 , 6143 ), and the SAT covers (0,300)(0,300)( 0 , 300 ). The final combined maps are then utilized for lensing B-mode template construction and parameter constraints. For comparative analysis, we additionally generate a LATS+SAT case by applying identical combination procedures to the NILC-cleaned maps from the LAT South (LATS) and SAT instruments.
The power spectra of selected polarization maps are shown in Fig. 5. It is evident that incorporating data from LATN (red dots) and LATS (blue dots) leads to a significant reduction in noise (see yellow dots), owing to the 20% sky overlap between the two experiments. Building on this, we further combine the data with SAT (see magenta plus signs). As expected, SAT contributes most of the signal at large scales ($\ell<300$), which is crucial for constraining $r$. It should be noted that for analyses incorporating SAT data (specifically the LATS+SAT and LATN+LATS+SAT combinations), we restrict the displayed data to multipoles $\ell<280$. Beyond this threshold, as the SAT contribution becomes negligible, the power spectra are expected to align with those derived from the LATS and LATN+LATS datasets respectively.
Figure 5: The power spectra of the selected polarization maps, both before and after map coaddition, are shown. The black solid lines represent the theoretical predictions. The ’Noise’ labels indicate contributions from both instrumental noise residuals and foreground residuals after NILC cleaning, while the ’Noisy’ labels refer to the NILC-cleaned maps. Labels with multiple experiments in parentheses correspond to the coadded maps, as described in Section 3.3. Note that for cases involving SAT, data are shown only for $\ell<280$, where the SAT contribution dominates. Additionally, the LATN and LATS do not fully overlap due to differences in their sky coverage, which is relevant for the Galactic foregrounds.
3.4 Lensing Reconstruction
3.4.1 Internal Lensing Reconstruction
We use the combined maps from the two ground-based LAT observations (LATN+LATS) to perform lensing reconstruction using the Quadratic Estimator (QE) method, following the implementation described in [28, 40]. To evaluate the improvement gained by incorporating LATN data, we also reconstruct the lensing potential using only the LATS maps.
For polarization reconstruction, we use a multipole range of (200,5600)(200,5600)( 200 , 5600 ), excluding multipoles below 200 to mitigate biases introduced in the delensing procedure. These biases arise from the overlap between the B-modes used for reconstructing $\phi$ and the B-modes intended for delensing (i.e., the observed B-modes), as noted by [53, 54]. For temperature reconstruction, we restrict the analysis to multipoles in the range (200,5200)(200,5200)( 200 , 5200 ) to relieve contamination from large-scale atmospheric noise and residual small-scale extragalactic foreground contributions, as discussed in [55, 56].
We implement the reconstruction using the Python package plancklens[40]. The spin-1 real-space (unnormalized) lensing deflection QE is given by:
${{}_{1}}\hat{d}(\hat{\mathbf{n}})=-\sum_{s=0,\pm 2}{{}_{-s}}\bar{X}(\hat{\mathbf{n}})\big{[}\sharp_{s}X^{\text{WF}}\big{]}(\hat{\mathbf{n}}).$ (3.11)
Here, $X=\{\Theta,E,B\}$ from observation, $\sharp$ denotes the spin-raising operator, which is associated with the gradient of a field [27], and the subscript $s$ represents the spin of the fields. The estimator takes as input the gradient of the Wiener-filtered maps, ${}_{s}X^{\text{WF}}(\hat{\mathbf{n}})$, and the real-space inverse-variance filtered maps, ${}_{s}\bar{X}(\hat{\mathbf{n}})$.
For filtering, we apply the optimal filtering method described in [57], implemented using the multi-grid preconditioned conjugate gradient algorithm. The iteration tolerance level $\epsilon$ is set to $10^{-5}$ for temperature filtering and $10^{-4}$ for polarization filtering, which is considered sufficiently precise for accurate lensing reconstruction. The root-mean-square (RMS) map required by this method is chosen to be an isotropic noise RMS map (we do not incorporate the hit-count variations in this work), with effective noise levels given in Table 3. These noise levels are determined from the power spectra of the null signal simulation, $X^{\text{NILC,null},i}$, for each experiment, thereby accounting for contributions from both instrumental noise and foreground residuals:
$\hat{\mathbf{n}}_{\text{lev}}=\sqrt{\langle\hat{C}_{\ell}^{\text{noise}}\rangle_{\ell}}\frac{60\cdot 180}{\pi},$ (3.12)
where $\langle\rangle_{\ell}$ denotes the average over the multipole range $2200<\ell<2600$ for temperature and $700<\ell<1000$ for polarization. This range is not strictly defined and is somewhat empirical, but it is chosen to suppress noisy modes at the smallest scales, where lensing reconstruction primarily relies on CMB information, as suggested in [29]. Additionally, we apply an apodized mask to reduce the creation of spurious gradients at the mask boundary [58] with an apodization scale of $6^{\prime}$, which has been shown to significantly accelerate the convergence of the iterative filtering process, particularly for temperature filtering.
Table 3: The effective white noise level used for lensing reconstruction filtering includes contributions from both instrumental noise and foreground residuals. Notably, in the overlapping regions, the polarization noise decreases by nearly $1/\sqrt{2}$ as expected, whereas the temperature noise shows only a slight reduction. We attribute this mainly to the large-scale foreground residuals, which are highly correlated since only one single foreground realization was used in the mock data.
nlev ($\mu$K-arcmin) LATN exclusive regions LATS exclusive regions Overlapping regions
Temperature 16.7 16.8 14.6
Polarization 8.0 8.0 5.7
Using the following relation, we derive a raw estimator for the lensing potential:
${}_{\pm 1}\hat{d}(\hat{\mathbf{n}})\equiv\left(\sum_{LM}\frac{\hat{g}_{LM}\pm i\hat{c}_{LM}}{\sqrt{L(L+1)}}\right){{}_{\pm 1}Y_{LM}(\hat{\mathbf{n}})},$ (3.13)
where the gradient mode $\hat{g}_{LM}$ is the quantity of interest, while the curl mode $\hat{c}_{LM}$ is expected to be negligible.
In practice, non-ideal factors such as the survey boundary and inhomogeneous noise introduce biases in the reconstruction. In this paper, we estimate the mean-field biases by averaging over simulation realizations and subsequently subtracting the estimated bias from the raw QEs. The final estimator is given by:
$\hat{\phi}_{LM}=\frac{1}{\mathcal{R}_{L}^{\phi}}\left(\hat{g}_{LM}-\langle\hat{g}_{LM}\rangle_{MC}\right),$ (3.14)
[…]
3.4.2 External Lensing Reconstruction
Under current CMB experimental conditions, internal lensing reconstruction typically suffers from a low signal-to-noise ratio (S/N) on small angular scales, primarily due to the dominance of reconstruction noise (notably the $N^{0}$ bias). To overcome this limitation, the incorporation of external large-scale structure (LSS) tracers—such as the Cosmic Infrared Background (CIB) and galaxy number density—has emerged as a promising strategy for enhancing the S/N of lensing reconstruction, especially at small scales. When properly filtered, these tracers can effectively complement internal delensing by providing additional information on the lensing potential at high multipoles [30, 31, 32].
Although the dominant contribution to large-scale lensed B-modes originates from lensing at intermediate angular scales ($L\sim\mathcal{O}(100)$; see Fig.2 of [32]), numerous studies have demonstrated that the inclusion of external tracers can nonetheless lead to a significant improvement in delensing efficiency [59, 16, 60, 61, 15].
To quantify the improvement in delensing performance achieved by incorporating large-scale structure (LSS) tracer data, we combine the internal lensing reconstruction from LATS and LATN+LATS with simulated external tracers—specifically the GNILC (Generalized NILC) CIB map from Planck PR2 [62, 63] and galaxy number density measurements from the Euclid survey [33]. These combinations are hereafter referred to as LATS+SAT+External and LATN+LATS+SAT+External, respectively.
The simulation of large-scale structure (LSS) tracer maps is based on theoretical models for their power spectra, which primarily comprise two components: (1) the underlying matter distribution, incorporating appropriate tracer biases, and (2) a noise term that accounts for experimental uncertainties, mainly arising from shot noise. We adopt the same configuration as used in our previous work [64]. A detailed description of the LSS models and the simulation setup can be found therein. For further background and reviews, see also [65, 66, 16, 14, 67, 34, 59, 33].
For the CIB, we use only the simulated maps at 353 GHz, where the GNILC CIB map has been shown to exhibit the highest correlation with CMB lensing [59]. Moreover, given the high signal-to-noise ratio (S/N) of the CIB auto-power spectrum over the relevant angular scales, the contribution of instrumental noise is negligible. Consequently, we omit instrumental noise in our theoretical modeling of the CIB.
For galaxy number density, we adopt the configuration of the Euclid survey, which achieves an effective galaxy number density of approximately 30 arcmin-2 [33]. To fully exploit the redshift information provided by Euclid, we divide the galaxy sample into five tomographic redshift bins, each serving as a tracer of the underlying matter distribution. It is worth noting that we assume the Euclid survey fully overlaps with the sky coverage of the ground-LATs’ configuration. In practice, some deviation is expected due to the differing scan strategies and scientific objectives of the two surveys, we leave a more detailed study to future work.
The tracers are combined through a linear combination of $I_{i}$:
$I=\sum_{i}c^{i}I_{i},$ (3.15)
where $I_{i}$ denotes the MV (minimum-variance) quadratic estimator (Eq. (3.14)), the CIB, and the galaxy number density from Euclid. The coefficients $c^{i}$ are determined by maximizing the correlation coefficient between the combined tracer and the true lensing convergence $\kappa$; the explicit form of $c^{i}$ can be found in [31]. It is important to note that the resulting combined tracer is no longer an unbiased estimator of the true convergence—instead, it effectively corresponds to a Wiener-filtered version. This approach is optimal for delensing, as it maximizes the overall delensing efficiency. [^3] Here, ”optimal” does not account for iterative lensing reconstruction [28, 29], and refers only to the combination of the quadratic estimator (QE) and LSS tracers.
3.5 Lensing B-mode Template Construction
3.5.1 Constructing LT on map-level
In this paper, we construct the lensing B-mode template using two delensing methods, as discussed in Section 2. To assess the improvement gained by incorporating LATN data and LSS tracers, we construct the lensing $B$-mode template (LT) separately for four cases: LATS+SAT, LATS+SAT+External, LATN+LATS+SAT, and LATN+LATS+SAT+External. In each case, the LT is constructed only within the overlapping region between the ground-based LATs and SAT. This restriction is not only due to the lack of lensing potential estimates in the non-overlapping regions, but also because of the absence of smaller-scale ($\ell>300$) lensed $E$-mode observations (see FIG.2 of [32]), both of which are essential for accurate LT construction.
The polarized maps for all cases are therefore masked with apodized masks to exclude sky regions beyond the LAT coverage and mitigate complications arising from sharp mask edges during the delensing procedure. This exclusion effectively prevents issues related to missing information at $\ell>300$ in these regions, as discussed in Section 3.3. The reconstructed lensing potential used for delensing is selected over the multipole range $20<L<5000$, and is Wiener-filtered as described in Section 4.1.
Before performing the delensing operation, we first apply a Wiener filter to the combined observed QU maps to enhance the signal-to-noise ratio (S/N), given by:
$\begin{gathered}Q_{\ell m}\Rightarrow\frac{C_{\ell}^{EE}}{C_{\ell}^{EE}+N_{\ell}^{EE}}Q_{\ell m},\\ U_{\ell m}\Rightarrow\frac{C_{\ell}^{EE}}{C_{\ell}^{EE}+N_{\ell}^{EE}}U_{\ell m},\\ \end{gathered}$ (3.16)
where the combined E-mode noise power spectrum is given by the weighted combination of each experiments:
$N_{\ell}^{EE}=\sum_{i}\omega_{i,\ell}^{2}N_{i,\ell}^{EE}.$ (3.17)
Regarding the Gradient-order lensing template method, we proceed by converting the spin-2 fields $Q+iU$ to spin-1 and spin-3 fields, and transform the spin-0 lensing potential field to spin-1-1- 1 and spin-1 fields using CMBlensplus [68]. Following this conversion, we compute the gradient lensing template according to Section 2. The resulting QU represents the lensing effect on E-mode, from which the lensing B-mode can be separated using harmonic transformation and EB leakage is corrected with an ”E-mode Recycling” method described in [49].
Regarding the Inverse-lensing method, we proceed by estimating the inverse deflection angle $\mathbf{d^{inv}}$ from the filtered lensing potential map $\hat{\phi}^{\text{MV}}$ and from combined tracer. This is done by iteratively solving Eq.(2.7) using CMBlensplus. Subsequently, we remap the filtered observed QU maps to obtain the delensed QU maps. These are then subtracted from the observed QU maps to derive the noisy lensing template QU maps. The lensing template B map can be separated through harmonic transformation from QU fields, and EB leakage should also be corrected.
3.5.2 Bandpower calculation and bias estimation
We compute the auto- and cross-bandpowers of lensing B-mode template (LT) and foreground-cleaned observed B-mode ($C_{b}^{\text{LT}}$ and $C_{b}^{\text{LT}\times\text{obs}}$) using the NaMaster[50] code. The multipole range for likelihood analysis (Section 3.6) spans from $\ell_{\text{min}}=20$ to $\ell_{\text{max}}=200$, with a binning width of $\Delta_{\ell}=20$.
As for the corresponding bias terms analyzed in Section 2.2, we input the required CMB maps and the potential maps to the same pipeline as described above to simulate all the bias terms, their auto- and cross-bandpowers are then also calculated using the NaMaster code. As mentioned in Section 2.2, we can also replace the reconstructed $\phi$ with the true $\phi$ in the baseline, which will cause some of the cross terms in the results to vanish. This allows us to isolate these terms by taking the difference between the baseline simulation results and these specific simulation results. We found that these two debiasing methods have similar effects, allowing us to choose either one flexibly. The average over 500 simulations of these terms serves as an estimate of the bias terms in the likelihood.
Note that the resulting lensing B-mode template (LT) is a filtered version of the lensed B-mode during to the suppression on signal from Wiener filter. So after debiasing, the auto- and cross-bandpowers are both filtered version and we then further compute the transfer function by signal-only simulation to compensate for the suppression on power-spectrum level. The transfer function for $C_{b}^{\text{LT}}$ is calculated as the ratio of the average power of the lensing B-mode template from signal-only simulation to the average lensed BB power. Besides, the transfer function for $C_{b}^{\text{LT}\times\text{obs}}$ is similarly calculated by replacing the numerator with the cross-spectrum between the constructed signal-only template and the signal lensed B-modes. We use these transfer functions to rescale $C_{b}^{\text{LT}}$ and $C_{b}^{\text{LT}\times\text{obs}}$ when fitting parameters.
3.6 Likelihood analysis and $r$ constraint
We investigate the parameter space using Cobaya [69, 70], which employs a Markov Chain Monte Carlo (MCMC) approach. To improve constraints on the tensor-to-scalar ratio $r$, rather than subtracting the lensing B-mode template (LT) from the observation (map-level delensing), we treat it as an additional pseudo-channel and compute the auto- and cross-bandpowers to construct the theoretical model (cross-spectral method delensing). As noted by [15], both methods yield comparable uncertainties. Our map channels consist of $B^{\text{NILC,obs}}$ and $B^{\text{LT}}$. The dataset vector is then defined as
$\mathbf{\hat{X}}_{b}=\left[C_{b}^{\text{NILC},\text{obs}},C_{b}^{\text{LT}\times\text{obs}},C_{b}^{\text{LT}}\right],$ (3.18)
which includes two auto-bandpowers and one cross-bandpower. We parameterize each element of $\mathbf{\hat{X}}_{b}$ using a two-parameter model:
$C_{b}^{\text{NILC,obs}}=rC_{b}^{\text{tens}}|_{r=1}+A_{L}C_{b}^{\text{lens}}+N_{b}^{BB,\text{NILC}},$ (3.19)
$C_{b}^{\text{LT}\times\text{obs}}={A_{L}}C_{b}^{\text{lens}},$ (3.20)
$C_{b}^{\text{LT}}=A_{L}C_{b}^{\text{lens}}+N_{b}^{\text{temp}},$ (3.21)
where the subscript $b$ denotes the bandpower bin, the parameter $A_{L}$ scales the intensity of the lensing B-mode. $N_{b}^{BB,\text{NILC}}$ represents the NILC residual power spectrum obtained from the null signal simulation $B^{\text{NILC,null}}$ (see Section 3.2). And $N_{b}^{\text{temp}}$ is the bias term as we discussed in Section 2.2.
The priors on the parameters are imposed as uniform distributions, as listed in Table 4. We assume a Gaussian likelihood to obtain the posterior distribution of parameters, which is a reasonable approximation given the large sky patch used in our analysis. The Gaussian approximation for the log-likelihood function adopted in this work is given by:
$\begin{gathered}-2\ \text{ln}\ \mathcal{L}(C_{b}|\hat{C_{b}})=(\mathbf{X}_{b}-\mathbf{\hat{X}}_{b})^{T}\mathbf{M_{b}^{-1}}(\mathbf{X}_{b}-\mathbf{\hat{X}}_{b})+\text{ln}\ |\mathbf{M_{b}}|,\end{gathered}$ (3.22)
where the vector $\mathbf{\hat{X}}_{b}$ contains the band-power estimates derived from the observed cut-sky maps, computed using the NaMaster code.
The covariance matrix $\mathbf{M_{b}}$ used in the likelihood analysis is derived from a baseline of 500 simulations, accounting for the effects of masking, cosmic variance, noise, foreground residuals, and other factors. The details of the simulation, foreground cleaning, and map combination procedures for the mock maps are provided in Section 3. Delensing is applied to these simulations to generate the lensing B-mode templates. Subsequently, all power spectra contained in $\mathbf{\hat{X}}_{b}$ are computed for each simulation using the NaMaster code, as described in Section 3.5. The covariance matrix is then estimated from these 500 realizations of $\mathbf{\hat{X}}_{b}$, with its elements given by:
$\mathbf{M}_{b,ij}=\langle\mathbf{\hat{X}}_{b,i}\mathbf{\hat{X}}_{b,j}\rangle-\langle\mathbf{\hat{X}}_{b,i}\rangle\langle\mathbf{\hat{X}}_{b,j}\rangle,$ (3.23)
where the average is taken over simulations, and the subscripts $i,j$ denote the elements of $\mathbf{\hat{X}}_{b}$.
4 Delensing Pipeline Performance
We now report the principal results from our delensing framework, organized into three key aspects: first, the reconstruction fidelity of the lensing potential quantified through both internal and external estimators; second, the effectiveness of lensing template generation in capturing B-mode contamination; and third, the cosmological parameter constraints derived from the delensed spectra. These results collectively establish the pipeline’s capability to reduce lensing-induced variance while preserving primordial signal, ultimately enabling tighter limits on $r$ compared to raw (non-delensed) measurements.
4.1 Results of lensing reconstruction
The final internal reconstructed lensing potential, $\hat{\phi}^{\text{MV}}_{LM}$, is shown in Figure 6. To highlight the lensing structures, we plot the Wiener-filtered deflection angle amplitude, defined as
$\hat{\alpha}^{\text{WF}}_{LM}=\sqrt{L(L+1)}\frac{C^{\phi\phi,\text{fid}}_{L}}{C^{\phi\phi,\text{fid}}_{L}+N^{(0),\text{ana}}_{L}}\hat{\phi}^{\text{MV}}_{LM}.$
The left panel presents the input data, while the middle panel shows the reconstructed deflection angle. The right panel displays their difference. It is evident that the reconstructed deflection successfully captures most of the features of the input data, and the homogeneous distribution of brightness and darkness in the difference map confirms the effectiveness of the reconstruction.
Figure 6: Reconstructed lensing map for LATN + LATS case. The left panel is the input lensing map; the middle panel is the reconstructed map from the MV quadratic estimator, the right panel is the difference between them. In order to highlight the lens structure, we plot the Wiener -filtered lensing deflection angle amplitude $\hat{\alpha}^{\text{WF}}_{LM}=\sqrt{L(L+1)}\hat{\phi}_{LM}^{\text{MV}}C^{\phi\phi,\text{fid}}_{L}/[C^{\phi\phi,\text{fid}}_{L}+N^{(0),\text{ana}}_{L}]$.
The power spectrum estimate of the reconstructed lensing convergence from a single simulation for both the LATS and LATN+LATS cases are shown in Fig. 7. These spectra have been binned and rescaled with a multiplicative correction following [40]. We observe a significant reduction in the standard deviation on large scales for the LATN+LATS case, primarily due to the increased sky coverage enabled by the addition of LATN. Furthermore, a notable reduction in reconstruction noise (dashed lines) is evident for the LATN+LATS case, attributed to the suppression of noise and foreground residuals resulting from the overlap between LATN and LATS.
Figure 8 shows the correlation coefficient $\rho=C_{L}^{\kappa\times I}/\sqrt{C_{L}^{\kappa\kappa}C_{L}^{II}}$ between each tracer and the true convergence field. Our results indicate that combining external tracers such as CIB and Euclid with internal CMB lensing reconstruction does not significantly enhance the correlation on large scales, where internal reconstruction already achieves a high signal-to-noise ratio (S/N). However, as expected, the external tracers substantially improve the correlation on smaller angular scales. This scale-dependent enhancement is particularly valuable, as it can contribute meaningfully to improving delensing performance under current experimental configurations.
Figure 7: The lensing power spectra from the internal reconstruction (MV) of a single simulation for both the LATS and LATN+LATS cases are shown. The light-blue and gold shaded regions represent the $1\sigma$ standard deviation estimated from 300 simulations for the LATS and LATN+LATS cases, respectively. The terms ”$N^{0}_{L}$” denote the semi-analytical leading-order reconstruction noise, which is used for Wiener filtering.
Figure 8: The correlation coefficient $\rho$ of each tracer, as well as the combined tracer (black), with the true convergence is shown. It is worth noting that improvements in $\rho$ at different angular scales contribute unequally to the effectiveness of delensing. In particular, the inclusion of external tracers significantly enhances the correlation on smaller scales compared to internal lensing reconstruction alone, which can lead to a non-negligible improvement in delensing performance.
4.2 Results of LT construction
We plot one of the delensed B-mode maps from our simulation set for $\ell<200$ in Figure 9. The delensed B-mode map is obtained by subtracting the gradient-order lensing B-mode template from the observation. After removing the noise component from the noisy lensing template at the map level, we divide the result by the square root of the transfer function, as calculated above, to compensate for signal suppression caused by filtering. It is evident that both the LATN+LATS+SAT and LATS+SAT cases mitigate part of the lensed B-modes. However, the former allows a larger fraction of the sky to be delensed, thanks to the additional data from LATN.
[…]
The power spectra of the NILC-cleaned B-modes and the lensing B-mode template (LT) are shown in Fig. 10. Compared to the LATS+SAT case, the power spectrum of the NILC-cleaned B-modes in the LATN+LATS+SAT case exhibits a slight reduction, attributed to its overlap with the other two experiments. Moreover, we observe a notable reduction in the standard deviation of the lensing B-mode template power spectrum in the LATN+LATS+SAT case. This improvement is primarily attributed to the increased sky coverage available for delensing, which enhances the statistical power of the reconstructed lensing signal. The reduction in standard deviation directly translates into a tighter constraint on the tensor-to-scalar ratio ($r$), as will be demonstrated in Section 4.3. Furthermore, the inclusion of LSS tracers (LATN+LATS+SAT+External case) leads to an additional reduction in the standard deviation by providing a more accurate estimate of the lensing potential.
Figure 10: The power spectra of the NILC-cleaned B-modes and the lensing B-mode template (LT) are shown for two cases, with all spectra properly debiased. The LT was generated using the gradient-order method, and the power spectra have been debiased and rescaled using the transfer functions. The black lines represent the theoretical predictions based on our models with best-fit parameters. In the top panel, the blue and green shaded areas indicate the $1\sigma$ standard deviation computed from 500 observation simulations for the LATN+LATS+SAT(same as LATN+LATS+SAT+External) and LATS+SAT(same as LATS + SAT + External) cases, respectively. In the lower two panels, the red, yellow and blue shaded areas represent the $1\sigma$ standard deviation of the LT for the LATN+LATS+SAT+External, LATN+LATS+SAT and LATS+SAT cases, respectively. As observed, all mean spectra in the lower two panels, after debiasing, align well with the theoretical predictions, which are precisely overlaid by the other curves. Moreover, a clear reduction in uncertainty is evident when comparing the LT derived from the LATN+LATS+SAT case to that from the LATS+SAT case.
4.3 Results on the $r$ constraint
The covariance matrices of the data vector $\mathbf{\hat{X}}_{b}$ for the LATS+SAT and LATN+LATS+SAT cases are shown in Fig. 11. A notable reduction in covariance is observed when comparing the LATN+LATS+SAT case to the LATS+SAT case, with the former exhibiting bluer overall. This reduction is primarily attributed to the increased sky coverage available for delensing.
[…]
Parameter Input value Prior
$r$ 0 $\mathcal{U}(-0.2,0.2)$
$A_{L}$ 1.000 $\mathcal{U}(0,1.2)$
(a) Covariance Matrix for LATS+SAT
(b) Covariance Matrix for LATN+LATS+SAT
Figure 11: The covariance matrix of the data vector $\mathbf{\hat{X}}_{b}$ for the LATS+SAT and LATN+LATS+SAT cases. Notice: The covariance values are shown in logarithmic scale for clarity. The nine slices represent the covariance matrices from bin 1 to bin 9, corresponding to the multipole range from $\ell=20$ to $\ell=200$, which are used for parameter constraints. We find an overall reduction in the matrix value for the LATN+LATS+SAT case (right panel).
We present the posterior distributions for the parameters using the two-parameter model with simulated data in Fig. 12. The degeneracy between $r$ and $A_{L}$ is progressively alleviated with the inclusion of the lensing B-mode template, as shown in the contour plots. Additionally, we observe a significant reduction in the uncertainty of $r$ with the inclusion of LATN data, regardless of whether the delensing procedure is performed.
The summarized results from the Gradient-order method and Inverse-lensing method are presented in Table 5 and Table 6, respectively. For the LATS+SAT case, the uncertainty in $r$ decreases from approximately $0.72\times 10^{-3}$ to $0.57\times 10^{-3}$ after incorporating the LT as an independent observation channel into the likelihood. This corresponds to a reduction of about 21%. Furthermore, we find that incorporating LATN data leads to a further reduction in the uncertainty of $r$. Even without delensing, this inclusion yields nearly an 8% improvement. Remarkably, when the lensing template is constructed with LATN data included, the uncertainty in $r$ is reduced by an additional 19%—a level of improvement comparable to that achieved by incorporating external LSS tracers ($\sim$ 17%). Once LATN is included, the marginal gain from adding external data results in a further reduction of approximately 10% in the uncertainty. We also find a strong consistency between the two delensing methods for most cases, further reinforcing the robustness of our results.
Figure 12: Posterior distributions of the baseline model parameters are shown for the LATS+SAT and LATN+LATS+SAT cases, with delensing and without delensing. The delensing is performed with the gradient-order method. The Gaussian approximation was used when performing the likelihood analysis. The lensing template (LT) was constructed using the gradient-order method. It is evident that the uncertainty in $r$ is reduced by including the LATN data in the likelihood analysis, for both ”With delensing” and ”Without delensing” cases.
Table 5: The mean and $1\sigma$ standard deviation of each parameter using Gradient-order method, with the LATS+SAT case, LATS+SAT+External case, LATN+LATS+SAT case and LATN+LATS+SAT+External case. Gaussian approximation likelihood was used in likelihood analysis. Notice that in the main text, the terms ”Without delensing” and ”With delensing” are used interchangeably with ”Before adding LT” and ”After adding LT” shown here, respectively.
Gradient-order Method Parameter Input value Before adding LT (LATS+SAT) Before adding LT (LATN+LATS+SAT) After adding LT (LATS+SAT) After adding LT (LATS+SAT+External) After adding LT (LATN+LATS+SAT) After adding LT (LATN+LATS+SAT+External) $r(\times 10^{3})$ 0 0.002±0.7220.002$\pm$ 0.7220.002 ± 0.722 0.004±0.666-0.004$\pm$ 0.666- 0.004 ± 0.666 0.009±0.5680.009$\pm$ 0.5680.009 ± 0.568 0.002±0.4410.002$\pm$ 0.4410.002 ± 0.441 0.002±0.4270.002$\pm$ 0.4270.002 ± 0.427 0.002±0.3540.002$\pm$ 0.3540.002 ± 0.354 $A_{L}$ 1 1.000±0.0181.000$\pm$ $0$ 0.018 1.000±0.0171.000$\pm$ $0$ 0.017 1.000±0.0151.000$\pm$ $0$ 0.015 1.000±0.0131.000$\pm$ $0$ 0.013 1.000±0.0141.000$\pm$ $0$ 0.014 1.000±0.0121.000$\pm$ $0$ 0.012
Table 6: The mean and $1\sigma$ standard deviation of each parameter using Inverse-lensing method, with the LATS+SAT case and LATN+LATS+SAT case. Gaussian approximation likelihood was used in likelihood analysis. Notice that in the main text, the terms ”Without delensing” and ”With delensing” are used interchangeably with ”Before adding LT” and ”After adding LT” shown here, respectively. We do not present cases with +External since there is no reason to expect any evident difference in the Gradient-order method.
Inverse-lensing Method Parameter Input value Before adding LT (LATS+SAT) Before adding LT (LATN+LATS+SAT) After adding LT (LATS+SAT) After adding LT (LATN+LATS+SAT) $r(\times 10^{3})$ 0 0.002±0.7220.002$\pm$ 0.7220.002 ± 0.722 0.004±0.666-0.004$\pm$ 0.666- 0.004 ± 0.666 0.011±0.5710.011$\pm$ 0.5710.011 ± 0.571 0.007±0.4290.007$\pm$ 0.4290.007 ± 0.429 $A_{L}$ 1 1.000±0.0181.000$\pm$ $0$ 0.018 1.000±0.0171.000$\pm$ $0$ 0.017 1.000±0.0151.000$\pm$ $0$ 0.015 1.000±0.0141.000$\pm$ $0$ 0.014
4.4 Limitations of current forecasts
Although we have attempted to produce realistic forecasts to assess the improvement in constraining the tensor-to-scalar ratio $r$ from a ground-based large-aperture telescope in the Northern Hemisphere, one limitation of these forecasts is the bias introduced by non-Gaussian foregrounds. Such foreground residuals are known to bias internal lensing reconstruction by several percent [71, 72, 73, 74], due to their response to the quadratic estimator (QE). This bias propagates through the delensing procedure, ultimately affecting the inferred constraint on $r$ [56]. As the impact of these effects is expected to be sub-dominant for LAT surveys such as SO, we defer a detailed investigation to future work.
5 Conclusions
In this paper, we present a pipeline designed to improve full-sky constraints on $r$ through large-scale B-mode delensing for future ground-based and satellite CMB observations. As baseline, we consider a scenario (denoted as LATS+SAT), where a ground-based large-aperture telescope in the Southern Hemisphere (LATS) enables the detection of the lensing potential, while a satellite-based small-aperture telescope (SAT) facilitates the detection of large-scale polarization with a significantly low noise level. Furthermore, we find that incorporating an additional ground-based LAT in the Northern Hemisphere (denoted as LATN+LATS+SAT) provides two key advantages. On the one hand, the overlap between experiments helps reduce noise. On the other hand, the expanded sky coverage allows for a larger area where the lensing B-mode template can be constructed, thereby enhancing the effectiveness of B-mode delensing. These two factors together significantly improve the constraints on $r$, further reducing its uncertainty compared to the LATS+SAT case. We also incorporate the external LSS tracers to the delensing procedure as comparison.
We apply the NILC method to remove foreground components in the observations from both LATs and SAT, utilizing six frequency channels for LATs and fifteen for SAT. We observe a significant reduction in the power of the cleaned maps, owing to the method’s objective of minimizing the total variance. For polarization, we expect noise and foreground residuals to have minimal impact on the delensing procedure, as discussed in our prior work [75]. For temperature, it is well known that the ILC method is less effective in suppressing compact extragalactic foregrounds. In our analysis, we retain these residuals, as we expect them to have a negligible impact on our results. A more detailed investigation of this issue is left for future work.
We perform internal lensing reconstruction using simulated data from ground-based LATs, incorporating both temperature and polarization information. External LSS tracers such as CIB and galaxy number density are further included in the pipeline.
The lensing B-mode template (LT) is constructed using simulated data from the combined observations of the satellite mission (SAT) and ground-based LATs, employing two delensing approaches. For parameter estimation, we extend the likelihood function to incorporate all auto- and cross-bandpowers between the LT and the observed B-modes. Our results show that internal delensing using southern-hemisphere data alone (LATS+SAT) reduces the uncertainty in the tensor-to-scalar ratio $r$ by approximately 21% compared to the no-delensing case. Furthermore, the inclusion of northern LAT (LATN) enables full-sky internal delensing and yields an additional $\sim$19reduction in the uncertainty of $r$, comparable to the improvement from including external LSS tracers ($\sim$17%). Once LATN is incorporated, the marginal gain from LSS tracers diminishes to around 10%.
We therefore conclude that for future full-sky missions targeting primordial gravitational waves, achieving full-sky internal delensing critically depends on the sky coverage of ground-based LATs. For LATs with SO-like sensitivity, the inclusion of a northern LAT provides an improvement in constraints on $r$ comparable to that achieved by incorporating external LSS data. This contribution is expected to become even more significant for future, ultra-sensitive LATs when used for full-sky delensing in support of space-based full-sky PGWs detection.
Acknowledgments
We thank Siyu Li, Yongping Li, Yepeng Yan, Zi-Xuan Zhang, Qian Chen for useful discussion. This study is supported by the National Natural Science Foundation of China No.12403005, the National Key R&D Program of China No.2020YFC2201601. We acknowledge the use of many python packages: CAMB [39], healpy and HEALPix [^4] https://healpix.sourceforge.io/ [47, 48], Lenspyx [41, 42], CMBlensplus [68], Plancklens[40], PySM3 [43, 44] , NaMaster [50], Cobaya [69, 70] and pyccl [76].